OpenStudy (anonymous):

alright dudes i got problem with equation of a line functions i am up to solving for b (y=mx+b) points are (-3.5) and (2,8) y=3/5x+b now i am stuck on solving this part 5=3/5*(-3) +b and 8= 3/5*(2) + b answer is suppose to be 34/5=b not sure how to get there

Directrix (directrix):

y = (3/5) x + b b is the y-intercept which is where the graph cuts the y-axis. To get on the y-axis, the x coordinate of the y intercept has to be zero. Think of the y-intercept as being in the form ( 0, b). You were given two points at the beginning. They are points on the line. So, they have to make the equation true. .

Directrix (directrix):

I chose point (2, 8) where x = 2 and y = 8 y = (3/5) x + b 8 = (3/5) * 2 + b Solve this for b and you'll have the y-intercept. Then, place it in this equation for b: y = (3/5) x + b @Kittens4Mittens

Directrix (directrix):

Now, another way to do this using technology. Go to the online equation of a line calculator site at: http://www.webmath.com/equline1.html You'll see something like that pictured on the attachment. Place the point coordinates into the calculator slots. Click on the find the equation tab. A discussion of how to do the problem follows along with the equation in y = mx + b form.

OpenStudy (anonymous):

thank you man!

Directrix (directrix):

You are welcome.